Suppose you have two clients who need your services for two years. One agreed to pay you $40,000 one year from now and another $40,000 in two years while the other paid $35,000 after one year, but $55,000 after two years. Assuming an interest rate of 5%, which one has a higher present value? Round off to the nearest dollar.
Answer #1

Present value= A/(1+i)^n

Present value of first client = 40000/(1+0.05) + 40000/(1.05)^2

= 74376.42

Present value of second client= 35000/(1.05) + 55000/(1.05)^2

= 83219.95

Thus the second client has higher present value.